OpenStudy (anonymous):

How do you simplify 2√5 + √45

OpenStudy (studygurl14):

\[2\sqrt{5} +\sqrt{9 \times 5} \rightarrow 2\sqrt{5}+3\sqrt{5} = 5\sqrt{5}\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Wait, where do you get the 3√5 from?

OpenStudy (anonymous):

\[3\sqrt{5}\] from

OpenStudy (studygurl14):

because \[\sqrt{9} = 3\] So...\[\sqrt{9\times5} = 3\sqrt{5}\]